In a certain number of two digits, the sum of the digits increased by 3, is twice the unit digit. If the number is divided by the sum of the digits, the quotient is 4 and the remainder is 3. Find the number.
Answer by mananth(16946)

let units digit be x
and tens digit be y
x+y+3 = 2x
x-y=3........(1)
(10y+x)= (x+y) *4 +3

subtract the equations
6y-3x =3
2y-x=1........(2)
add 1 & 2
y=4
therefore x=7
Number is 47